OpenStudy (anonymous):

A class is having a contest to see who can guess the number of marbles in a jar. A guess that is 10 or less from the actual number of marbles in the jar wins a school T-shirt. The jar has 486 marbles. Which of these inequalities best represents a guess that wins a school T-shirt?

OpenStudy (mathmate):

@ludvic Do you have any suggested answer?

OpenStudy (anonymous):

@mathmate a)\[-10\le x \le10\] b)\[\x-486|\ge 10\] c)\[\x-486|\le 10\] d)\[x-486 \le 10\]

OpenStudy (joannablackwelder):

Can you eliminate any answers that just don't make sense?

OpenStudy (mathmate):

You can also draw a number line to help you: 

OpenStudy (mathmate):

Which of the answers do you choose and why?